# The effect of interactions on elastic cavitation

### 20 Mar 2026 v1

@@ Abstract
In this study, we characterize the effect of cavity interaction on cavitation pressure for (i) a nearby rigid interface and (ii) a neighboring identical cavity. For cavities near a rigid interface, our analysis shows that the cavitation pressure increases as the initial cavity-interface distance decreases, starting from the bulk value $p_{\rm c,bulk}$ for a distant cavity and approaching the cavitation pressure value for a defect situated at an interface ($p_{\rm c,int}\approx 3.5\mu$) as the cavity approaches the interface boundary. In contrast, interacting cavities exhibit a non-monotonic dependence of the cavitation pressure on the initial inter-cavity distance $d$: the threshold approaches the bulk value $p_{\rm c,bulk}$ for distant cavities and reaches a maximum of $p_{c}\approx 2.8\mu$ at $d\approx 5.7R$, where $R$ is the initial cavity radius. These results provide, for the first time, a quantitative estimate of cavitation pressure for these two interaction-effect extensions of the classical cavitation problem and will help guide modeling of cavitation-driven processes in porous, biological, and multi-material solids.
@@ 2 Results
Proximity to an interface or another cavity (Fig. 1) affects the cavitation pressure relative to its bulk value. To quantitatively describe these effects, we introduce the following notation. The dimensionless pressure is defined as $\overline{p}=p/\mu$, where $p$ is the remotely applied pressure and $\mu$ is the shear modulus of the material. The dimensionless volume ratio is defined as $\overline{V}=V/V_{0}$, where $V_{0}$ is the initial, undeformed cavity volume, and $V$ is its deformed volume. The critical value of pressure $p$ which causes unstable cavity expansion is the cavitation pressure $p_{c}$, and $\overline{p}_{c}=p_{c}/\mu$ is the dimensionless cavitation pressure. As a reference value, the bulk cavitation pressure for a spherical cavity in an infinite, incompressible neo-Hookean medium is $\overline{p}_{c,\rm bulk}=2.5$. The cavitation pressure for expansion of a defect located exactly at the interface between the elastic body and a bonded rigid substrate was established in a prior study [16] as $\overline{p}_{\rm c,int}\approx 3.5$, also called the interfacial cavitation pressure. The key results of this study are as follows:
@@
Fig. 2 shows the pressure-volume relation for a sequence of distances (expressed in the dimensionless form $d/R$ where $R$ is the radius of the cavity and $d$ is the distance of its center from the rigid interface in the undeformed configuration) and Fig. 2 shows the corresponding post-processed cavitation pressures. The cavitation pressure increases monotonically as the undeformed cavity location gets closer to the interface, with peak values approaching $\overline{p}_{\rm c,int}$—a significant increase from the bulk cavitation pressure (which is the asymptotic limit of the cavitation pressure for $d/R\to\infty$). As the distance between the cavity and the interface increases, the pressure-volume curve approaches that of the bulk cavitation case (dashed black curve in Fig. 2, defined in eq. (3)).
@@
The cavitation pressure varies non-monotonically with the undeformed inter-cavity half-distance $d$, as shown in Fig. 3. A peak is observed at $d/R\approx 5.7$ where $\overline{p}_{c}\approx 2.8$, highlighting a maximal cavity interaction effect; for large $d/R$, the response approaches the single-cavity limit $\overline{p}_{c}=2.5$ (dashed black curve in Fig. 3).
@@ 4.1 Discretization and Simulation Setup
We carry out the simulations using the open source finite element software FEniCS [1, 30, 31]. Incompressibility is enforced via a mixed displacement-pressure ($u-\tilde{p}$) formulation using the lowest-order Taylor-Hood element pair (quadratic elements for displacement and linear elements for pressure), with a perturbed Lagrangian treatment of the volumetric constraint ($J-1=\tilde{p}/K$, where $\tilde{p}$ is the Lagrange multiplier interpreted as a hydrostatic pressure and $K=10^{12}\times G$ is a penalty parameter with the interpretation of a bulk modulus). The weakform setup in FEniCS parallels the implementation in [42] (see also [2]). The domain dimensions used in the simulation are shown in Figure 5. The mesh is especially refined near the cavity to ensure sufficient accuracy in capturing the large deformations in the vicinity of the cavity. Figure 6 shows an example of the meshes employed; here, the minimum element size is $R/100$, and 46,968 triangular elements are used to discretize the domain using the software Gmsh [15]. Mesh convergence study is presented in A.
@@ 4.2 Cavitation pressure through curve extrapolation
Given that the simulations show a monotonic increase in cavity volume with applied pressure for all cases, the cavitation pressure can be determined by the asymptotic value of the pressure as $\overline{V}\to\infty$. While the simulations are able to achieve large values of $\overline{V}$, the pressure-volume curves need to be extrapolated to determine the asymptotic cavitation pressure value. We achieve this by employing a curve-fitting procedure for the pressure-volume curve and then using the fitted curve to estimate the cavitation pressure. Following testing of different basis choices for the curve-fitting, Laguerre polynomials were chosen since they yielded the most stable convergence behavior for cavitation pressure predictions, while reproducing the monotonicity/non-monotonicity behaviour of the simulated pressure-volume curves with respect to $d/R$ in the predicted cavitation pressures.
Noting that Laguerre polynomials of degree $m$ are given by
@@
the functional fit for the dimensionless pressure-volume curve is
@@
where $a_{m}$ are the fitting coefficients. Since $L_{m}(0)=1$, the cavitation pressure is given by the sum of the coefficients as
$\overline{p}_{c}=\lim_{\overline{V}\to\infty}\hat{p}=\sum_{m=0}^{N}a_{m}$ (6)
A satisfactory fit is achieved when the computed root-mean-square error (RMSE) $\leq 10^{-5}$ and the cavitation pressure stabilizes with $N$ (where $N+1$ is the number of fitting terms in eq. (5)). The RMSE is defined as
@@
We found that $N=40$ was sufficient to achieve a satisfactory fit across all $d/R$ values. The fitting procedure is illustrated for two different $d/R$ values, for the cavity-interface interaction problem in Fig. 7, and for the cavity-cavity interaction problem in Fig. 7.
@@
Figure 7: Fitting procedure results for: (a) cavity-cavity, $d/R=20$ (b) cavity-cavity, $d/R=1.1$ (c) cavity-interface, $d/R=20$ and (d) cavity-interface, $d/R=1.1$.
@@ 5 Conclusion
For a cavity near a rigid interface, the cavitation pressure decreases monotonically with increasing cavity–interface spacing, with values ranging between $p_{c}\approx 3.5\mu$ for small $d/R~(\approx 1)$ and bulk cavitation value of $p_{\rm c,bulk}=2.5\mu$ for large $d/R$. In cavity-cavity interaction case, the cavitation pressure is non-monotonic as a function of inter-cavity half-distance, with a peak cavitation pressure $p_{c}\approx 2.8\mu$ exhibited around $d/R\approx 5.7$. The noticeable enhancement effect in both cases, with respect to the classical bulk cavitation pressure, indicates the importance of taking into account the presence of interface and nearby cavities, which is in turn likely to affect the mechanical response and failure in a variety of engineering and biological systems.
@@ Appendix A Mesh Sensitivity Analysis
$V=\int_{X_{1}^{2}+(X_{2}-d)^{2}=R^{2}}\pi r^{2}\left(-n_{1}\right)\,\mathrm{d}s$ (10)
@@
Appendix B Fitting of Cavitation Pressure vs Inter-cavity Distance Curve
To obtain smooth representations of the cavitation pressure as a function of spacing, $d/R$, the discrete values shown as black dots in Figs. 2 and 3 are approximated using analytical smoothing functions. These fits provide continuous curves that capture the underlying trends of the numerical data. For the cavity-interface interaction case in Fig. 2, the behavior is captured using a sigmoidal error–function fit,
$\overline{p}_{c}(x)=L+\frac{U-L}{2}\left[1+\operatorname{erf}\left(\frac{x_{0}-x}{w}\right)\right],\qquad x=d/R,$ (14)
where $L=3.6$, $U=2.5$, $x_{0}=5.35$ and $w=3.15$. For the cavity-cavity interaction case in Fig. 3, the data were approximated using a rational polynomial,
$\overline{p}_{c}(x)=\frac{r_{1}x^{3}+r_{2}x^{2}+r_{3}x+r_{4}}{x^{3}+q_{1}x^{2}+q_{2}x+q_{3}},\qquad x=d/R,$ (15)
where $\{r_{1},r_{2},r_{3},r_{4}\}=\{2.522,-17.15,-33.26,413.2\}$ and $\{q_{1},q_{2},q_{3}\}=\{-6.494,-17.44,172.9\}$ are fitting coefficients.

### 30 Mar 2026 v2

@@ Abstract
In this study, we characterize the effect of cavity interaction on cavitation pressure for (i) a nearby rigid interface and (ii) a neighboring identical cavity. For cavities near a rigid interface, our analysis shows that the cavitation pressure increases as the initial cavity-interface distance decreases, starting from the bulk value $p_{\rm c,bulk}$ for a distant cavity and approaching the cavitation pressure value for a defect situated at an interface ($p_{\rm c,int}\approx 3.5\mu$) as the cavity approaches the interface boundary. In contrast, interacting cavities exhibit a non-monotonic dependence of the cavitation pressure on the initial inter-cavity distance $d$: the threshold approaches the bulk value $p_{\rm c,bulk}$ for distant cavities and reaches a maximum of $p_{\rm c}\approx 2.67\mu$ at $d\approx 3.8R$, where $R$ is the initial cavity radius. These results provide, for the first time, a quantitative estimate of cavitation pressure for these two interaction-effect extensions of the classical cavitation problem and will help guide modeling of cavitation-driven processes in porous, biological, and multi-material solids.
@@ 2 Results
Proximity to an interface or another cavity (Fig. 1) affects the cavitation pressure relative to its bulk value. To quantitatively describe these effects, we introduce the following notation. The dimensionless pressure is defined as $\overline{p}=p/\mu$, where $p$ is the remotely applied pressure and $\mu$ is the shear modulus of the material. The dimensionless volume ratio is defined as $\overline{V}=V/V_{0}$, where $V_{0}$ is the initial, undeformed cavity volume, and $V$ is its deformed volume. The critical value of pressure $p$ which causes unstable cavity expansion is the cavitation pressure $p_{\rm c}$, and $\overline{p}_{\rm c}=p_{\rm c}/\mu$ is the dimensionless cavitation pressure. As a reference value, the bulk cavitation pressure for a spherical cavity in an infinite, incompressible neo-Hookean medium is $\overline{p}_{\rm c,\rm bulk}=2.5$. The cavitation pressure for expansion of a defect located exactly at the interface between the elastic body and a bonded rigid substrate was established in a prior study [16] as $\overline{p}_{\rm c,int}\approx 3.5$, also called the interfacial cavitation pressure. The key results of this study are as follows:
@@
Fig. 2 shows the pressure-volume relation for a sequence of distances (expressed in the dimensionless form $d/R$ where $R$ is the radius of the cavity and $d$ is the distance of its center from the rigid interface in the undeformed configuration) and Fig. 2 shows the corresponding post-processed cavitation pressures. The cavitation pressure increases monotonically as the undeformed cavity location gets closer to the interface, attaining $\overline{p}_{\rm c}\approx 3.44$ at d/R=1.1d/R=1.1—a significant increase from the bulk cavitation pressure (which is the asymptotic limit of the cavitation pressure for $d/R\to\infty$). As the distance between the cavity and the interface increases, the pressure-volume curve approaches that of the bulk cavitation case (dashed black curve in Fig. 2, defined in eq. (3)).
@@
The cavitation pressure varies non-monotonically with the undeformed inter-cavity half-distance $d$, as shown in Fig. 3. A peak is observed at $d/R\approx 3.8$ where $\overline{p}_{\rm c}\approx 2.67$, highlighting a maximal cavity interaction effect; for large $d/R$, the response approaches the single-cavity limit $\overline{p}_{\rm c}=2.5$ (dashed black curve in Fig. 3).
@@ 4.1 Discretization and Simulation Setup
We carry out the simulations using the open source finite element software FEniCS [1, 30, 31]. Incompressibility is enforced via a mixed displacement-pressure ($u-\tilde{p}$) formulation using the lowest-order Taylor-Hood element pair (quadratic elements for displacement and linear elements for pressure), with a perturbed Lagrangian treatment of the volumetric constraint ($J-1=\tilde{p}/K$, where $\tilde{p}$ is the Lagrange multiplier interpreted as a hydrostatic pressure and $K=10^{12}\times G$ is a penalty parameter with the interpretation of a bulk modulus). The weakform setup in FEniCS parallels the implementation in [42] (see also [2]). The domain dimensions used in the simulation are shown in Figure 5. The mesh is especially refined near the cavity to ensure sufficient accuracy in capturing the large deformations in the vicinity of the cavity. Figure 6 shows an example of the meshes employed; here, the minimum element size is $R/250$, and 168758 triangular elements are used to discretize the domain using the software Gmsh [15]. Mesh convergence study is presented in A.
@@ 4.2 Cavitation Pressure through Curve Extrapolation
Given that the simulations show a monotonic increase in cavity volume with applied pressure, where the pressure appears to asymptote in all cases, the cavitation pressure can be determined by the limiting value of the pressure as $\overline{V}\to\infty$. While the simulations are able to achieve large values of $\overline{V}$, the pressure-volume curves need to be extrapolated to determine the asymptotic cavitation pressure value. We achieve this by employing a curve-fitting procedure for the pressure-volume curve and then using the fitted curve to estimate the cavitation pressure.
Motivated by the expression for the pressure-volume relation for the bulk cavitation case in eq. (3), we employ polynomial basis functions in terms of the variable $\overline{V}^{-1/3}$ for the fitting. Specifically, we choose Laguerre polynomials of degree $m$, given by
@@
since they are orthonormal in $x\in[0,\infty)$, where the functional fit for the dimensionless pressure-volume curve is given as
@@
and $a_{m}$ are the fitting coefficients. Given that $L_{m}(0)=1$, the cavitation pressure is determined by the sum of the coefficients as
$\overline{p}_{\rm c}=\lim_{\overline{V}\to\infty}\hat{p}=\sum_{m=0}^{N}a_{m}$ (6)
The root-mean-square error (RMSE) of the fitted curve is defined as
@@
We find that the RMSE decreases with increasing number of fitting terms $N+1$, however for $N>5$, the condition number of the matrix associated with the linear equation system to solve for the least squares fitting of the parameters $a_{m}$ becomes poor, reflecting over-parameterization and poor numerical stability of the fitted coefficients. Consequently, we cap the value of $N$ at 5, which gives satisfactory fit (low RMSE) for all $d/R$ values and report the average of the cavitation pressures obtained using the fits for $N=4$ and $N=5$ (the choice of $N\geq 4$ covers all the terms in eq. (3)). Further, we note that using a regular polynomial basis instead of the Laguerre polynomial basis yields identical results reflecting the robustness of the fitted cavitation pressure. The fitting procedure is illustrated for two different $d/R$ values, for the cavity-interface interaction problem in Fig. 7-7, and for the cavity-cavity interaction problem in Fig. 7-7.
@@
Figure 7: Fitting procedure results for: (a) cavity-interface, $d/R=20$ (b) cavity-interface, $d/R=1.1$ (c) cavity-cavity, $d/R=20$ and (d) cavity-cavity, $d/R=1.1$.
@@ 5 Conclusion
For a cavity near a rigid interface, the cavitation pressure decreases monotonically with increasing cavity–interface spacing, with values ranging between $p_{\rm c}\approx 3.44\mu$ for small d/R(=1.1)d/R~(=1.1) and bulk cavitation value of $p_{\rm c,bulk}=2.5\mu$ for large $d/R$. In cavity-cavity interaction case, the cavitation pressure is non-monotonic as a function of inter-cavity half-distance, with a peak cavitation pressure $p_{\rm c}\approx 2.67\mu$ exhibited around $d/R\approx 3.8$. The noticeable enhancement effect in both cases, with respect to the classical bulk cavitation pressure, indicates the importance of taking into account the presence of interface and nearby cavities, which is in turn likely to affect the mechanical response and failure in a variety of engineering and biological systems.
@@ Appendix A Mesh Sensitivity Analysis
$V=\int_{X_{1}^{2}+(X_{2}-d)^{2}=R^{2}}\pi r^{2}\left(-n_{1}\right)\,\mathrm{d}s(\mathbf{X})$ (10)